# ALMAS: an Autonomous LLM-based Multi-Agent Software Engineering Framework

### v1

@@ II Related Work
In software engineering, Large Language Models (LLMs) are applied to a diverse array of tasks, including code summarization and comprehension [7, 8, 9, 10], program synthesis [11, 12], code translation [13, 14], automated program repair [15, 16, 17], and test generation [18, 19]. Many of these applications focus on the coding aspect of software development. However, studies indicate that only 15% to 35% of human effort is dedicated to the implementation phase of the SDLC [20]. Motivated by this observation and recent surveys that highlight both the promise and limitations of current approaches [21, 22], our work hypothesizes that integrating automated techniques across the entire SDLC can substantially improve development processes. Accordingly, our ALMAS framework spans multiple phases in a multi-agent setting.
@@
Recent surveys provide comprehensive overviews of LLM-based multi-agent systems in software engineering. For example, [21] outlines the opportunities and challenges of these systems, while [22] categorizes existing approaches based on workflow, infrastructure, and inherent challenges like context management and communication overhead. Although these surveys offer analyses, they do not address concrete strategies for mitigating failure modes such as task verification errors or inter-agent misalignment.
Cognitive Assistance and Architectural Approaches
Complementary to these surveys, recent work has focused on specialized aspects of automated assistance and system architecture. For instance, one study leverages cognitive models to enhance developer support by mimicking human memory processes [23], and another proposes a modular, distributed architecture that orchestrates task-specific agents [24]. While these efforts have advanced our understanding of cognitive support and scalable system design, they tend to address only isolated portions of the SDLC. In contrast, ALMAS integrates these insights by providing an end-to-end framework that not only covers agile role decomposition and dynamic agent orchestration but also specifically tackles issues related to context-window limitations and attention dilution.
Failure Analysis in Multi-Agent Frameworks
The study in [25] critically examines common failure modes in LLM-based multi-agent systems, including ineffective task verification and misalignment during inter-agent communication. This analysis underscores the need for designs that address such issues directly. ALMAS responds to these challenges through agile role alignment, dynamic code summarization, and a novel retrieval strategy that collectively mitigate these failure modes.
Code Augmentation and Retrieval Challenges.
Several approaches have augmented code generation by embedding retrieval mechanisms to extend contextual awareness. For instance, Agentless [26] employs file hierarchy representations for code localization and SWE-Agent [4] and OpenHands [27] iteratively retrieves code segments to support bug fixes. Although these methods improve targeted code fixes, they typically do not address broader challenges such as context-window limitations in large codebases. ALMAS builds on these techniques by integrating a novel retrieval strategy (Meta-RAG) with dynamically generated code summaries, extending its applicability to both bug fixing and new feature development [28].
In summary, while prior work in conversational code assistants and retrieval-based approaches [29, 30, 4, 26, 27] has advanced our understanding of interactive programming support and code augmentation, these studies largely address isolated aspects of the SDLC. Broad surveys [21, 22] provide valuable context yet fall short of offering integrated solutions. ALMAS unifies these strands by orchestrating distinct agents for sprint planning [31], code localization [28], generation, and review [32] [33], into a cohesive framework that addresses key challenges such as limited context windows, attention dilution, and inter-agent misalignment.
@@ III Vision
An overview of the vision for the ALMAS agents and components is presented in Figure 1, and described in more details throughout this section.
III-A Software Planning
The framework initiates with the Sprint Agent, which acts as both the Product Manager and Scrum Master. Previous research has shown that unclear and incomplete requirements can impose significant costs and risks during the development process [34, 35]. Therefore, upon receiving a high-level task from the user, the Sprint Agent evaluates the task for clarity and completeness, making the necessary improvements to ensure it is well-defined. The Sprint agent then devises a stepwise plan to achieve the task’s objectives, breaking it down into sub-tasks complemented by additional descriptions, acceptance criteria, and effort estimation. Acceptance criteria are crucial for generating unit tests and conducting code reviews. Meanwhile, effort estimation, a common practice in agile teams, is facilitated by the Sprint agent using few-shot examples of previous estimations to adopt the appropriate metric for estimation and improve accuracy as published in our previous work [31].
The Supervisor Agent also plays a pivotal role in routing the allocation of sub-tasks to the appropriate agents, optimizing for cost and performance by maintaining an inventory of various code LLMs with different sizes and expertise [36].
III-B Context-Aware Development
Context-aware development is facilitated by the Summary Agent and Control Agent, beginning with a preprocessing operation crucial for handling pre-existing code repositories. The Summary Agent condenses the codebase into structured natural-language replicas, tackling the context length issue with LLMs. Each code file is transformed into a summary file, containing a concise sentence that encapsulates the file’s objective and responsibility. Additionally, a hierarchical structure is built providing one-sentence summaries for each code unit, including file-level scripts, functions, classes, and class-level functions. These summaries serve as a programming-language-independent context for other agents, except for the Code Agents that directly interact with the code. This approach not only reduces the cost associated with using LLMs by minimizing token usage but also enhances their performance by allowing them to engage with the codebase in natural language—a modality more suited to most LLMs than code.
The Control Agent utilizes Retrieval Augmented Generation (RAG) over summaries to localize changes required for each sub-task. It selects a list of code units needed to be provided as context to the code agent to guide the code generation process. Moreover, the Sprint agent uses summaries as context while breaking down the initial task. This part of the framework is referred to as Meta-RAG, evaluated in our previous work [28].
III-C Collaborative Development
The Developer Agent embodies a bundle of multiple agents that collaboratively fulfill the developer’s responsibilities. It receives sub-tasks and localized code units from the Supervisor Agent, retrieves the corresponding code from the codebase, and works to implement the required changes. The collaborative nature of the Developer Agent ensures that each sub-task is addressed efficiently, leveraging the strengths of different agents with different capabilities to achieve the desired outcomes. Importantly, the framework is designed to work in parallel with human developers, allowing for seamless integration within an agile team. Developers can choose which agents to incorporate into their workflow, thanks to the framework’s modular design. This adaptability means that a team might opt to utilize only specific agents, such as the Sprint and Peer Review agents, to complement their existing processes. This collaborative approach not only enhances productivity but also fosters a harmonious interaction between automated agents and human developers. Furthermore, this agent can be easily replaced with other state-of-the-art code generation solutions, such that the solutions can be ensembled for better outcomes.
III-D Cost Efficiency
ALMAS is designed with cost efficiency in mind, leveraging several strategies to reduce cost without sacrificing performance. By condensing the codebase into structured natural-language replicas and keeping it up to date with the interim changes, the framework reduces token usage in the long run, which is a significant factor in LLM cost [28]. The Supervisor Agent enhances cost efficiency by strategically routing tasks to the most suitable LLMs, considering their specialty, size, and cost [36]. By having access to a diverse set of LLMs, it can select the optimal option for each task, optimizing resource use. This modular design allows teams to customize the framework to their needs, further reducing costs through selective agent utilization.
@@ III-D Cost Efficiency
III-E Validation, Verification, and Error Handling
Validation checks are integral to the framework, ensuring the accuracy and reliability of the processes involved. Drawing from human parallels, the Developer Agent undertakes validation processes to check for formatting and compile issues in the generated code, ensuring adherence to required standards. It executes unit tests to validate the integrity of the code, ensuring that all previous functionality remains intact. In addition, the Peer Agent conducts a comprehensive review of the code changes, assessing functionality alignment, vulnerability checks [32], performance evaluation, hallucinations [33], formal verification [37], and code quality. This review culminates in a report attached to the pull request, providing recommendations for acceptance or rejection to the human developer.
The framework is also equipped with mechanisms to recover from unexpected issues that may arise during the process. Error handling during unit tests is a primary recovery strategy; if a test fails, the error log is returned to the Control Agent, which localizes the necessary changes to address the issue (repeating process from step 3 in Figure 1). The Supervisor Agent monitors the progress of all developer agents, maintaining a history of actions taken by each agent to ensure correct direction. In cases where the framework encounters issues that LLMs cannot resolve in a certain (tunable) number of tries, control may be handed over to a human developer, with a summarized history of actions providing insights into progress and challenges. This handover process ensures continuity and resilience in the development process.
III-F Tool Usage and Integration
The framework is designed to easily integrate with various tools used in the SDLC, specifically CI/CD tools, code versioning and task management tools, such that it fits seamlessly into the workflow of development teams. Furthermore, the framework can be integrated with IDEs such as VS Code or IntelliJ via plugins, allowing direct interaction with developers and other agile team members. These integrations ensure that the framework is not only a standalone solution but also a complementary tool that enhances the existing workflows of development teams.

### v2

@@ II Related Work
Large Language Models (LLMs) have been widely adopted in software engineering for tasks such as code summarization [7, 8], program synthesis [9], code translation [10], automated repair [11, 12], and test generation [13]. Despite this focus on coding, only a fraction of developer effort is spent on implementation [14]. Motivated by this and recent surveys highlighting both the potential and current gaps of LLM-based tools [15, 16], our work proposes that automation across the full SDLC can yield greater benefits. The ALMAS framework is designed to address this by supporting multiple SDLC phases in a multi-agent context.
@@
Recent surveys [15, 16] review LLM-driven multi-agent systems, discussing workflows, infrastructure, and challenges like context management. However, they stop short of proposing concrete solutions for issues such as task verification failures or agent misalignment.
Cognitive and Architectural Approaches
Other studies have explored cognitive models for developer support [17] and modular agent architectures [18], but these typically address only specific SDLC stages. ALMAS extends these ideas, offering an end-to-end framework that manages agile role decomposition, dynamic orchestration, and context limitations.
Failure Analysis in Multi-Agent Systems
Work by [19] analyzes common failures in LLM-based multi-agent systems, such as poor task verification and communication breakdowns. ALMAS addresses these with agile role alignment, dynamic summarization, and a novel retrieval strategy to reduce such failures.
Code Augmentation and Retrieval
Recent methods have enhanced code generation with retrieval mechanisms—Agentless [20], SWE-Agent [4], and OpenHands [21]—to improve bug fixing. Yet, these often overlook broader issues like context-window limits in large codebases. ALMAS advances this area by combining a new retrieval approach (Meta-RAG) with dynamic code summaries, supporting both bug fixes and feature development [22].
In summary, while prior work has advanced individual aspects of LLM-based software engineering, most solutions are fragmented. ALMAS unifies these advances, coordinating agents for planning [23], localization [22], generation, and review [24, 25], to address persistent challenges like limited context, attention dilution, and agent misalignment.
@@ III Vision
Software Planning: ALMAS begins with the Sprint Agent, acting as Product Manager and Scrum Master. Since unclear requirements can increase development risks [26, 27], the Sprint Agent refines user tasks for clarity and completeness, then breaks them into sub-tasks with descriptions, acceptance criteria, and effort estimates. Acceptance criteria support unit testing and code review, while effort estimation leverages few-shot learning from past examples [23]. The Supervisor Agent allocates sub-tasks to the most suitable LLMs, optimizing for cost and performance by maintaining a diverse agent pool [28].
Context-Aware Development: The Summary and Control Agents enable context-aware development by preprocessing code repositories. The Summary Agent generates concise, structured natural-language summaries for each code unit, addressing LLM context limitations and reducing token costs. These summaries provide language-agnostic context for other agents. The Control Agent uses Retrieval Augmented Generation (RAG) over these summaries to localize relevant code for each sub-task, a process called Meta-RAG we introduced in [22]. The Sprint Agent also uses these summaries for task decomposition.
Collaborative Development: The Developer Agent coordinates multiple specialized agents to implement sub-tasks, using localized code and summaries. This collaborative setup allows agents to work alongside human developers, who can selectively integrate agents like Sprint or Peer Review into their workflow. The modular design supports easy replacement or ensembling of code generation solutions, enhancing flexibility and productivity.
Cost Efficiency: ALMAS is designed with cost efficiency in mind, leveraging several strategies to reduce cost without sacrificing performance. By condensing the codebase into structured natural-language replicas and keeping it up to date with the interim changes, the framework reduces token usage in the long run, which is a significant factor in LLM cost [22]. The Supervisor Agent enhances cost efficiency by strategically routing tasks to the most suitable LLMs, considering their specialty, size, and cost [28]. By having access to a diverse set of LLMs, it can select the optimal option for each task, optimizing resource use. This modular design allows teams to customize the framework to their needs, further reducing costs through selective agent utilization.
Validation, Verification, and Error Handling: Validation is built in: the Developer Agent checks code formatting, compilation, and runs unit tests to ensure correctness. The Peer Agent reviews code for functionality, vulnerabilities [24], performance, hallucinations [25], formal verification [29], and quality, providing a report for human review. For error handling, failed tests trigger the Control Agent to localize and address issues, while the Supervisor Agent tracks agent actions. If automated recovery fails after several attempts, control passes to a human developer with a summarized action history, ensuring resilience.
Tool Usage and Integration: ALMAS integrates seamlessly with SDLC tools—CI/CD, version control, task management—and can connect to IDEs like VS Code or IntelliJ via plugins. This ensures it fits naturally into existing development workflows, serving as both a standalone and complementary solution.